OpenStudy (anonymous):

how will i evaluate log(base3/2) 16/81 ? anyone !! please help me

OpenStudy (yttrium):

is it \[\log_{\frac{ 3 }{ 2 }} \frac{ 16 }{ 81 } \]

OpenStudy (psymon):

So first off: \[\log_{b}M = x\]is the same exact thing as: \[b^{x} = M\] So with your problem, you have: \[\log_{3/2}\frac{16}{81}= x\]Meaning we can rewrite this as \[(\frac{ 3 }{ 2 })^{x} = \frac{16}{81}\] So we need to raise 3/2 to some power that gets us 16/81. So one way to visually help us out is another rewrite. (3/2)^x can be rewritten as: \[\frac{ 3^{x} }{ 2^{x} }\]Now we see something doesnt match. Powers of 3 go 3 9 27 81 and powers of 2 go 2 4 8 16. So from the looks of it, the powers of our numerators matchthe denominator of 16/81 and the powers of the denominator match the numerator of 16/81. In order to flip our fraction so numerators match numerators, we need to have a negative exponent. Negative exponents basically flip numbers from numerator to denominator and vice versa. So, as you may have notice, 3 9 27 81, 3^4 gives us 81 and 2 4 8 16, 2^4 gives us 16. But since we need it upside down, we just say that instead of the correct exponent being 4, we need it to be -4. So the answer in all of this is -4 because: \[(\frac{ 3 }{ 2 })^{-4} = \frac{16}{81}\], which also is the same as saying: \[\log_{3/2}\frac{16}{81} = -4\]

OpenStudy (yttrium):

if it is, then \[(\frac{ 3 }{ 2 })^x = \frac{ 16 }{ 81 }\] Therfore, \[(\frac{ 2 }{ 3 })^{-x} = \frac{ 16 }{ 81 }\] Hence, \[(\frac{ 2 }{ 3 })^{-4} = \frac{ 16 }{ 81 }\] So x = -4

OpenStudy (anonymous):

oh tnx... i got it... can i ask another question?